# A Survey of Reinforcement Learning For Economics

### 15 Mar 2026 v2

@@ 6.7 Simulation Study: Bellman Error and Value Error in Offline Policy Evaluation
Two algorithms are trained on $\mathcal{D}$.11111111150,000 gradient steps per seed, 3 seeds; two-layer MLP with 64 hidden units, Adam at $5\times 10^{-4}$; target network updated every 500 steps. BRM minimizes $\mathbb{E}_{\mathcal{D}}[(Q(s,a)-(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})))^{2}]$ where both $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ use the current network; gradients flow through both sides simultaneously. The key consequence is that the network can zero the residual at an observed pair $(s,a)$ by co-moving $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ together, rather than moving either toward $Q^{*}$. This is the opposite of supervised learning, where labels are fixed external targets that do not move with the model weights. FQE prevents this by using a frozen target network for $Q(s^{\prime},a^{\prime})$; the network must reduce $Q(s,a)$ toward a target that does not respond to its own gradient steps. Every 500 steps we record the Bellman error on $\mathcal{D}$ (current network on both sides) and the value error $\frac{1}{|\mathcal{F}|}\sum_{(s,a)\in\mathcal{F}}|Q_{\theta}(s,a)-Q^{*}(s,a)|$ over all 4,795 feasible pairs. As a supervised baseline, OLS regression of $\log c$ on $\log k$ and $\log z$ is estimated on expanding windows.112112112Noise $\sigma=0.30$; windows from $n=10$ to 2,0002{,}000; held-out test set of 500 transitions.
@@ 6.8 Discussion and Recommendations
7 Reinforcement Learning Algorithms: An Illustrated Example
The preceding chapters derived convergence rates, identified failure modes, and catalogued the tradeoffs governing reinforcement learning algorithms. This section runs nine algorithms on a common environment to observe these properties empirically, with particular attention to per-state value function convergence.
7.1 Setup
The environment is a $5\times 5$ deterministic gridworld with five actions (up, down, left, right, stay). Each non-terminal step incurs a penalty of 0.1-0.1; reaching the goal in the bottom-right corner yields +10+10. A small state-dependent perturbation ($\epsilon=0.001$) breaks the symmetry between equally-long paths, ensuring a unique optimal policy.113113113Without the perturbation, any shortest path from start to goal is optimal. The perturbation subtracts $\epsilon\cdot i/N^{2}$ from each step reward, where $i$ is the destination state index. Since moving down increases the state index by $N$ while moving right increases it by 1, the agent strictly prefers right over down at every interior state. The discount factor is $\gamma=0.95$ and each episode is capped at 50 steps. The optimal path requires 8 steps, yielding an optimal discounted return of 9.30.114114114Shared hyperparameters across all learning algorithms: learning rate $\alpha=0.1$ (decay 0.99995/episode) for value-based methods, $\alpha=0.01$ for policy gradient methods; $\epsilon$-greedy exploration from 1.0 to 0.01 (decay 0.9995/episode); trace decay $\lambda=0.9$; NPG actor step size $\eta=0.2$, critic step size 0.5 (both decay 0.99995/episode). DQN uses a two-layer network with 64 hidden units, replay buffer of 50,000 transitions, batch size 64, learning rate $10^{-3}$, target network update every 100 episodes, and $\epsilon$-decay 0.9999. PPO uses clip ratio 0.2, 4 optimization epochs per batch, GAE $\lambda=0.95$, learning rate $0.01$ (decay 0.99995/episode). All learning algorithms run for 500,000 episodes with seed 42.
With only 25 states, the value function $V^{*}(s)$ can be displayed in full. The grid ranges from $V^{*}(s_{0})=6.38$ at the top-left corner to $V^{*}=10.0$ at the state adjacent to the goal, with values increasing monotonically along the optimal path. This small scale makes it feasible to track convergence at every state individually, rather than relying on aggregate metrics that can mask important differences between algorithms.
Two planning methods serve as baselines: value iteration (VI) and policy iteration (PI), both requiring the full transition model $P(s^{\prime}|s,a)$. Seven model-free learning algorithms observe only sampled transitions: Q-learning, SARSA, Q($\lambda$), DQN, REINFORCE with value baseline, Natural Policy Gradient (NPG), and Proximal Policy Optimization (PPO). DQN replaces Q-learning’s lookup table with a two-layer neural network (64 hidden units) as function approximator, training on minibatches sampled from a replay buffer. Algorithm definitions appear in Sections 4.15.5; this section focuses on empirical behavior.
Table 8 reports results for all nine algorithms.
Table 8: Performance of nine algorithms on the $5\times 5$ gridworld. “Ep to $V^{*}$ conv” is the first checkpoint at which $\max_{s}|V(s)-V^{*}(s)|<0.1$ across all 25 states; “—” indicates the algorithm never reaches this threshold. All seven learning algorithms achieve the optimal return of 9.30, but only Q-learning, Q($\lambda$), and DQN converge to $V^{*}$ everywhere.
Algorithm Return Agreement (%) $V^{*}$ RMSE Max $|V{-}V^{*}|$ Ep to $V(s_{0})$ Ep to $V^{*}$ conv Time (s)
Planning (full model)
VI 9.30 100.0 0.00 0.00 9 iter 0.001
PI 9.30 100.0 0.00 0.00 9 iter 0.118
Temporal difference
Q-Learning 9.30 100.0 0.00 0.00 340 610 226.9
SARSA 9.30 88.0 1.67 5.53 7600 693.2
Multi-step
Q($\lambda$) 9.30 100.0 0.00 0.00 340 550 218.1
Deep RL
DQN 9.30 92.0 0.01 0.01 1600 1600 2781.3
Policy gradient
REINFORCE 9.30 76.0 2.20 5.71 132000 142.7
NPG 9.30 44.0 4.52 9.72 200 118.3
PPO 9.30 52.0 1.88 5.42 30000 258.5
7.2 Planning Versus Learning
VI converges in 9 iterations (0.001 seconds). On this deterministic gridworld, the iteration count matches the MDP diameter: the longest shortest path from any state to the goal is 8 steps, and each Bellman backup propagates value one step outward from the terminal state.115115115The $\gamma^{k}$ contraction bound from Section 5.1 predicts roughly $\log(\epsilon^{-1})/\log(\gamma^{-1})\approx 360$ iterations for $\epsilon=10^{-8}$, making it loose by a factor of 40 on this structured problem.
PI converges in 9 policy-improvement steps (0.018 seconds). At each step, PI solves the linear system $V^{\pi}=r^{\pi}+\gamma P^{\pi}V^{\pi}$ and propagates value information across the full state space simultaneously, rather than one step at a time as in VI.
The fastest learning method, Q($\lambda$), needs 20 episodes to reach 99% of the optimal return; NPG and PPO need 50; SARSA needs 80. The gap between 0.001 seconds of planning and 20–80 episodes of learning is the price of not knowing the transition model $P(s^{\prime}|s,a)$.
7.3 Value Function Convergence
All seven learning algorithms achieve the optimal return of 9.30, yet only three converge to $V^{*}(s)$ for all 25 states (Table 8). This distinction, invisible in the return metric, is the central finding of this comparison.
Q-learning and Q($\lambda$) are off-policy and learn $Q^{*}$ regardless of the behavior policy. Both achieve $\max_{s}|V(s)-V^{*}(s)|<0.1$ by episode 610 and 550 respectively, with final RMSE below 0.01. Q($\lambda$) converges slightly faster due to eligibility traces propagating the terminal reward backward through the full 8-step trajectory within a single episode.
DQN is also off-policy and converges to $V^{*}$ at all 25 states by episode 1600, with RMSE 0.006 and $\max_{s}|V(s)-V^{*}(s)|=0.01$. These errors are comparable to the tabular methods. The cost is wall-clock time: DQN requires 2921 seconds versus 25 seconds for tabular Q-learning, a $118\times$ slowdown from neural network training overhead on a problem small enough for exact tabular representation.
SARSA, REINFORCE, NPG, and PPO are on-policy and learn the value of the exploration policy $\pi_{\epsilon}$, not $\pi^{*}$. SARSA leaves 9 of 25 states unconverged at episode 500,000, with a maximum error of 5.53 concentrated at the state farthest from the goal. REINFORCE leaves 15 of 25 states unconverged (max error 5.71), NPG leaves 12 of 25 unconverged (max error 9.72), and PPO leaves 16 of 25 unconverged (max error 5.42).
Table 9 tracks $\max_{s}|V(s)-V^{*}(s)|$ at five episode checkpoints. For Q-learning and Q($\lambda$), the maximum error drops below 0.1 by episode 1,000 and continues falling. DQN follows the same trajectory with a slight delay. The on-policy methods show a different pattern: SARSA’s maximum error remains above 5.0 throughout training, and the policy gradient methods plateau at similar levels, confirming that the error is structural rather than a matter of insufficient training.
Table 9: Maximum per-state value error $\max_{s}|V(s)-V^{*}(s)|$ over episodes (rolling average over $\pm 3$ snapshots). Off-policy methods converge uniformly; on-policy methods plateau at persistent errors driven by rarely-visited states.
Algorithm Ep 100 Ep 1K Ep 10K Ep 100K Ep 500K
Q-Learning 7.64 0.00 0.00 0.00 0.00
SARSA 8.94 5.90 5.53 5.53 5.53
Q($\lambda$) 4.64 0.00 0.00 0.00 0.00
DQN 8.40 1.17 0.03 0.02 0.01
REINFORCE 8.24 6.94 6.85 6.47 5.71
NPG 9.72 9.72 9.72 9.72 9.72
PPO 8.33 7.12 6.60 6.22 5.42
Table 10 tracks the fraction of states at which each algorithm’s greedy policy matches $\pi^{*}$. Q-learning and Q($\lambda$) reach 100% policy agreement by episode 10,000. DQN achieves 92% at its best checkpoints but oscillates substantially between snapshots, a consequence of neural network weight updates shifting the argmax at states with similar Q-values. SARSA stabilizes at 88% despite never converging in value. REINFORCE, NPG, and PPO plateau at 44–76%.
Table 10: Policy agreement $(\%$ of states where $\pi(s)=\pi^{*}(s))$ over episodes (rolling average over $\pm 3$ snapshots). Tabular off-policy methods reach full agreement. DQN’s average is depressed by inter-snapshot oscillation. Policy gradient methods plateau below 80%.
Algorithm Ep 100 Ep 1K Ep 10K Ep 100K Ep 500K
Q-Learning 59 91 100 100 100
SARSA 74 82 88 88 88
Q($\lambda$) 48 94 100 100 100
DQN 52 70 61 61 56
REINFORCE 72 76 76 76 76
NPG 48 48 48 44 44
PPO 49 52 52 52 52
7.4 Optimal Returns Without Optimal Value Functions
NPG reaches 99% of optimal return at episode 50 but never converges to $V^{*}$ at 12 of 25 states. The per-state error pattern reveals why: NPG’s value function is accurate along the optimal path (states near the diagonal from start to goal) but wildly inaccurate at off-path states. At state (1,4), NPG’s error is 9.72, nearly the full magnitude of $V^{*}$. The agent has learned to act optimally on the equilibrium path without learning what off-equilibrium states are worth.
This is the distinction between on-path and off-path behavior. An agent can achieve optimal performance in its observed environment, acting correctly on the equilibrium path, without having solved the complete decision problem for every contingency. In structural estimation terms, the agent’s observed behavior reveals optimal choices, but its counterfactual predictions (what would it do at states it never visits?) are unreliable. Off-policy methods sidestep this problem by learning $Q^{*}$ from any data, including exploratory transitions; on-policy methods learn the value of the policy they actually follow, which is the exploration policy during training.
Figure 13 makes this visible. At $5\times 5$ resolution, each state’s error is individually readable. Q-learning shows a wavefront of decreasing error spreading from the goal; by the final episode, the entire grid is white (error near zero). NPG, REINFORCE, and PPO show persistent dark patches at states the optimal policy rarely visits, with errors exceeding 5.0 at states in the far corners.
Figure 13: Learned value function $V(s)$ at percentage-based checkpoints of each algorithm’s $V^{*}$-convergence time $T$, defined as the first episode where $\max_{s}|V(s)-V^{*}(s)|<0.1$; algorithms that never converge use T=500, [^0] T=500{,}000. All columns share the same color scale, so a converged algorithm’s grid visually matches the $V^{*}$ reference column. Off-policy methods (Q-learning, Q($\lambda$), DQN) converge to $V^{*}$ everywhere. On-policy methods (SARSA, REINFORCE, NPG, PPO) show persistent discrepancies at off-path states even at the final episode.
Figure 14 shows the corresponding policy grids. Q-learning’s arrows match $\pi^{*}$ everywhere by the final checkpoint. NPG and PPO show incorrect arrows concentrated at the same off-path states where value errors are largest, since the agent has never needed to act at these states, so its policy there reflects the softmax initialization rather than learned behavior.
Figure 14: Policy grids at percentage-based checkpoints of $V^{*}$-convergence time $T$ (same definition as Figure 13). Arrows indicate the greedy action at each state; dots indicate the stay action. Off-policy methods converge to $\pi^{*}$ everywhere. On-policy methods retain incorrect actions at states far from the optimal path.
@@ 8 Reinforcement Learning for Optimal Control
A handful of organizations have deployed reinforcement learning beyond simulation, achieving measurable gains on specific large-scale problems. Each deployment required substantial domain engineering and scientific tuning; these remain exceptional cases rather than standard practice. I review the most prominent field deployments, including ride-hailing dispatch at DiDi, data center cooling at Google, hotel revenue management, and financial order execution, before concluding with a simulation study on the bus engine replacement problem.
@@ 8.1 Ride-Hailing Dispatch
Qin et al. (2021) formalized DiDi’s order dispatching as a semi-Markov decision process116116116A semi-Markov decision process generalizes the standard MDP by allowing variable time between decisions. The discount factor $\gamma^{\tau_{k}}$ depends on the actual time elapsed $\tau_{k}$ rather than applying a fixed per-step discount. where each driver is an independent agent. The state of a driver consists of location (discretized into hexagonal zones) and time (bucketed into intervals). The action is the order assigned to the driver, with the option to remain idle. The reward is the trip fare. State transitions are determined by trip destinations, as completing an order transports the driver from origin to destination, changing their spatial state. The stochasticity arises from future demand, which determines the available actions at each state. The per-driver value function $V^{\pi}(s)$ represents the expected cumulative fare a driver can earn from state $s$ under dispatching policy $\pi$:
@@
where $\hat{r}_{i}$ is the predicted fare for order $i$, $\hat{\tau}_{i}$ is the estimated trip duration, and $s_{j}^{\prime}$ is the destination state. The value function is learned via temporal-difference methods from historical trip data. DiDi’s Cerebellar Value Network (CVNet; Tang et al., 2019) uses hierarchical coarse-coding117117117Coarse-coding represents a value function as a weighted sum over overlapping rectangular or hexagonal tiles that partition the state space (Sutton and Barto, 2018); each state activates the tiles that contain it, and a hierarchical variant stacks tiles at multiple resolutions to allow both coarse and fine generalization simultaneously. with a multi-resolution hexagonal grid, enabling transfer learning118118118Transfer learning initializes a model for a new task (a new city) with parameters trained on related tasks (existing cities), on the assumption that learned representations of traffic and demand patterns generalize across contexts and require less data to reach good performance in the new setting. across cities and robustness to data sparsity in low-traffic zones.
@@
Han et al. (2022) reported complementary results from Lyft, where the dispatching system optimizes driver assignment and repositioning jointly. Their value decomposition architecture119119119Value decomposition decomposes the platform’s global matching objective into individual driver value functions, each estimable via temporal-difference learning. The global optimum is recovered by optimizing the sum of individual values. assigns credit to individual driver decisions within the global objective. The production system demonstrated improvements in rider wait times and driver utilization, providing a second data point suggesting that similar approaches may transfer across platforms.
@@
8.2 Data Center Cooling
Data centers consume enormous amounts of energy, with cooling systems accounting for 30–40% of total power usage. Google applied machine learning to autonomous HVAC control, achieving substantial efficiency gains while maintaining temperature safety constraints.
Gao (2014) reported initial results from deploying neural network predictions to guide data center operations. The system used over 100 sensor readings, including temperatures at multiple locations, coolant flow rates, power draw, and external weather conditions, to predict Power Usage Effectiveness (PUE).120120120PUE is the ratio of total facility energy to IT equipment energy. A PUE of 1.0 means all energy goes to computing; the industry average is approximately 1.6. Operators used these predictions to adjust equipment setpoints. This assisted approach achieved a 40% reduction in cooling energy and a 15% improvement in overall PUE.
Lazic et al. (2018) subsequently deployed a fully autonomous system using model-predictive control121121121Model-predictive control (MPC) solves a finite-horizon optimal control problem at each step using a predictive model of system dynamics, applies the first action, then re-solves at the next step with the updated state; the receding-horizon structure makes it robust to model error because the plan is continually revised. with learned dynamics. The state space consists of temperatures, flow rates, and equipment states across multiple zones. Actions are setpoints for chillers, fans, and pumps. The system learns a neural network dynamics model $\hat{f}$ from operational data:
$s_{t+1}=\hat{f}(s_{t},a_{t})+\epsilon_{t}.$ (61)
At each control interval (5 minutes), the system optimizes over a horizon of future states using the learned model, subject to temperature constraints.
$\min_{a_{0},\ldots,a_{H}}\sum_{h=0}^{H}c_{\text{energy}}(a_{h})\quad\text{s.t.}\quad T_{\min}\leq T_{\text{zone},i}(\hat{s}_{h})\leq T_{\max}\;\forall i,h.$ (62)
The MPC approach provides interpretability and safety guarantees that pure model-free RL cannot. Actions are projected onto the feasible set defined by temperature constraints, with uncertainty estimates from the learned model used to maintain conservative margins. Human operators can override at any time. Table 12 summarizes the reported performance.
Table 12: Google data center cooling results from Gao (2014) and Lazic et al. (2018).
Metric Improvement Method
Cooling energy -40% Neural network predictions
PUE -15% Assisted operation
Additional efficiency -12% Autonomous MPC
Temperature violations 0 Safety constraints
The system operates continuously across multiple data centers, processing millions of sensor readings per day and controlling thousands of setpoints. Importantly, this is model-predictive control with learned dynamics models and persistent human oversight, not autonomous model-free RL; the learned model adapts online to seasonal variations, equipment degradation, and workload changes, but operators retain override authority at all times.
@@ 8.3 Hotel Revenue Management
Hotels face a dynamic pricing problem with perishable inventory, since rooms unsold by check-in yield zero revenue. The state $(t,I,d)$ consists of days until check-in, remaining rooms, and a demand intensity indicator. Actions are discrete price levels. Booking arrivals follow a Poisson process with intensity depending on demand conditions; each arrival books with probability decreasing in price. The Bellman equation is:
$V_{t}(I,d)=\max_{p\in\mathcal{A}}\mathbb{E}\left[p\cdot B+V_{t-1}(I-B,d^{\prime})\mid I,d,p\right],$ (63)
where $B$ is the number of bookings at price $p$ and $d^{\prime}$ is the next-period demand state.
Gallego and van Ryzin (1994) established the analytical benchmark for dynamic pricing with Poisson demand. Under their intensity control framework, the optimal price trajectory can be characterized when demand follows specific functional forms. The Gallego-van Ryzin solution provides a theoretical upper bound, and any practical pricing system should approach this performance in settings where the model assumptions hold.
Chen et al. (2023) applied deep Q-learning to hotel revenue management, training on simulated booking data under parameterized demand models. The DQN agent uses state features (normalized time, normalized inventory, demand indicators) and outputs Q-values for each price level. Table 13 compares performance against baselines on held-out test periods.
Table 13: Hotel revenue management results from Chen et al. (2023).
Method Mean Revenue vs. Fixed Price
Fixed price 2,847
Myopic pricing 2,962 +4.0%
EMSR-b heuristic 3,156 +10.9%
Exact DP 3,241 +13.8%
DQN 3,198 +12.3%
The DQN agent achieves 98.7% of the DP optimal revenue while scaling to capacities where exact DP is infeasible.122122122EMSR-b (Expected Marginal Seat Revenue, variant b) is the standard heuristic in airline and hotel revenue management. It computes booking limits by comparing the expected marginal revenue of reserving a unit for a higher fare class against selling it now. Training time is nearly constant in hotel capacity, whereas DP computation grows linearly. The learned policies exhibit interpretable patterns, with prices increasing as inventory depletes and as check-in approaches, with larger adjustments during high-demand periods.

### 17 Mar 2026 v3

@@ 6.7 Simulation Study: Bellman Error and Value Error in Offline Policy Evaluation
Two algorithms are trained on $\mathcal{D}$.11911911950,000 gradient steps per seed, 3 seeds; two-layer MLP with 64 hidden units, Adam at $5\times 10^{-4}$; target network updated every 500 steps. BRM minimizes $\mathbb{E}_{\mathcal{D}}[(Q(s,a)-(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})))^{2}]$ where both $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ use the current network; gradients flow through both sides simultaneously. The key consequence is that the network can zero the residual at an observed pair $(s,a)$ by co-moving $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ together, rather than moving either toward $Q^{*}$. This is the opposite of supervised learning, where labels are fixed external targets that do not move with the model weights. FQE prevents this by using a frozen target network for $Q(s^{\prime},a^{\prime})$; the network must reduce $Q(s,a)$ toward a target that does not respond to its own gradient steps. Every 500 steps we record the Bellman error on $\mathcal{D}$ (current network on both sides) and the value error $\frac{1}{|\mathcal{F}|}\sum_{(s,a)\in\mathcal{F}}|Q_{\theta}(s,a)-Q^{*}(s,a)|$ over all 4,795 feasible pairs. As a supervised baseline, OLS regression of $\log c$ on $\log k$ and $\log z$ is estimated on expanding windows.120120120Noise $\sigma=0.30$; windows from $n=10$ to 2,0002{,}000; held-out test set of 500 transitions.
@@ 7 Reinforcement Learning for Optimal Control
A handful of organizations have deployed reinforcement learning beyond simulation, achieving measurable gains on specific large-scale problems. Each deployment required substantial domain engineering and scientific tuning; these remain exceptional cases rather than standard practice. I review the most prominent field deployments, including ride-hailing dispatch at DiDi, data center cooling at Google, hotel revenue management, and financial order execution, before concluding with a simulation study on the bus engine replacement problem. In each case, the RL agent’s parameters were updated during a training phase conducted in simulation or on historical data, and the resulting policy was deployed with fixed weights (Section 2).
@@ 7.1 Ride-Hailing Dispatch
Qin et al. (2021) formalized DiDi’s order dispatching as a semi-Markov decision process121121121A semi-Markov decision process generalizes the standard MDP by allowing variable time between decisions. The discount factor $\gamma^{\tau_{k}}$ depends on the actual time elapsed $\tau_{k}$ rather than applying a fixed per-step discount. where each driver is an independent agent. The state of a driver consists of location (discretized into hexagonal zones) and time (bucketed into intervals). The action is the order assigned to the driver, with the option to remain idle. The reward is the trip fare. State transitions are determined by trip destinations, as completing an order transports the driver from origin to destination, changing their spatial state. The stochasticity arises from future demand, which determines the available actions at each state. The per-driver value function $V^{\pi}(s)$ represents the expected cumulative fare a driver can earn from state $s$ under dispatching policy $\pi$:
@@
where $\hat{r}_{i}$ is the predicted fare for order $i$, $\hat{\tau}_{i}$ is the estimated trip duration, and $s_{j}^{\prime}$ is the destination state. The value function is learned via temporal-difference methods from historical trip data. DiDi’s Cerebellar Value Network (CVNet; Tang et al., 2019) uses hierarchical coarse-coding122122122Coarse-coding represents a value function as a weighted sum over overlapping rectangular or hexagonal tiles that partition the state space (Sutton and Barto, 2018); each state activates the tiles that contain it, and a hierarchical variant stacks tiles at multiple resolutions to allow both coarse and fine generalization simultaneously. with a multi-resolution hexagonal grid, enabling transfer learning123123123Transfer learning initializes a model for a new task (a new city) with parameters trained on related tasks (existing cities), on the assumption that learned representations of traffic and demand patterns generalize across contexts and require less data to reach good performance in the new setting. across cities and robustness to data sparsity in low-traffic zones.
@@
Han et al. (2022) reported complementary results from Lyft, where the dispatching system optimizes driver assignment and repositioning jointly. Their value decomposition architecture124124124Value decomposition decomposes the platform’s global matching objective into individual driver value functions, each estimable via temporal-difference learning. The global optimum is recovered by optimizing the sum of individual values. assigns credit to individual driver decisions within the global objective. The production system demonstrated improvements in rider wait times and driver utilization, providing a second data point suggesting that similar approaches may transfer across platforms.
@@ 7.2 Hotel Revenue Management
Budget hotel chains face a capacity allocation problem: how to dynamically distribute rooms across rate segments defined by discount level, with booking channels such as direct platforms and online travel agencies mapped to segments offering discounts ranging from less than 15% to over 40%. Demand is uncertain, cancellations are difficult to forecast, and hotel managers resist black-box optimization systems that override their judgment. Chen et al. (2023) deployed reinforcement learning at China Lodging Group (CLG), a budget hotel chain operating approximately 2,000 hotels across China, and conducted field experiments measuring the impact of RL-based capacity allocation on actual hotel revenue.
Their system uses a two-step design. The RL agent observes the state $(t,s)$, where $t$ indexes the booking period within a $T=10$-period episode and $s$ is the average revenue per room sold to date, and selects an average discount level $a\in\{10\%,20\%,30\%\}$. A linear program then converts this scalar recommendation into a feasible capacity allocation across rate segments, accounting for each hotel’s channel preferences.125125125The RL component uses a modified on-policy Monte Carlo method with $\varepsilon$-greedy exploration, updating Q-values from realized returns after each completed episode. Training on real operational data rather than simulated environments means the method adapts to demand non-stationarity without requiring an explicit demand model. This decomposition addresses practitioner acceptance, since managers understand a single discount recommendation, and circumvents the need to explicitly model demand arrivals or cancellation rates.
The field experiment randomly assigned five Hanting-brand hotels in Shanghai to the treatment group from ten candidates, with 271 additional Shanghai hotels serving as a donor pool for synthetic control estimation.126126126Synthetic control constructs a weighted combination of untreated hotels whose pre-treatment trend matches each treated hotel, avoiding the selection bias of simple before-after comparisons. Table 10 reports the average treatment effects over the pilot period (March–June 2015).
Table 10: Field experiment results from Chen et al. (2023), measured via synthetic control.
Metric Improvement $p$-value
Occupancy rate (OR) $5.15\%$ 0.1361
Average daily rate (ADR) $5.93\%$ 0.1160
Revenue per available room (RevPAR) $11.80\%$ 0.0090
The RevPAR gain is heterogeneous across treatment hotels; some improved primarily via higher occupancy rates, others via higher average daily rates, and others via both channels simultaneously.
7.3 E-Commerce Dynamic Pricing
E-commerce platforms face a pricing problem at a scale that defeats human specialists. Alibaba’s Tmall.com lists millions of SKUs across thousands of product categories, each requiring daily price adjustments that account for demand elasticity, competitor behavior, inventory levels, and promotional calendars. Liu et al. (2019) deployed deep reinforcement learning agents for automated pricing on Tmall.com beginning July 2018, conducting field experiments on thousands of SKUs over several months.
The pricing MDP for product $i$ has state $s_{i,t}\in\mathbb{R}^{m}$ comprising four feature groups: price features (current and historical prices, price-to-cost ratio), sales features (units sold, conversion rate), customer traffic features (unique visitors $\text{uv}_{i,t}$, page views), and competitiveness features (price rank among similar products). The pricing period is $d=1$ day. Actions are either discrete, $a_{i,t}\in\{1,\ldots,K\}$ indexing price bins uniformly spaced between product-specific bounds $P_{i,\min},P_{i,\max}]$, or continuous, $a_{i,t}\in\mathbb{R}$. The reward is the difference of revenue conversion rates (DRCR):
$r_{i,t}=\frac{\text{revenue}_{i,t}}{\text{uv}_{i,t}}-\frac{\text{revenue}_{i,t-\tau}}{\text{uv}_{i,t-\tau}},$ (61)
where $\text{uv}_{i,t}$ is unique visitors in period $t$ and $\tau$ is a reference lag.127127127DRCR normalizes revenue by traffic to remove demand fluctuations unrelated to pricing. The differencing further removes product-specific level effects, yielding a reward signal that is more concave than raw revenue and improves convergence stability.
Two algorithms are deployed: DQN for the discrete action formulation and DDPG for continuous actions. Both are pre-trained from demonstrations using historical specialist pricing actions (DQfD, DDPGfD) to address cold-start.128128128Standard A/B testing is infeasible because Chinese e-commerce regulations prohibit displaying different prices to different customers for the same product simultaneously. Liu et al. (2019) instead evaluate using difference-in-differences against “simi-products” (similar products not subject to algorithmic pricing) as controls.
Table 11: Field experiment results from Liu et al. (2019) on Tmall.com. DRCR improvement is relative to the simi-product control group.
Experiment Method DRCR Improvement Duration
Markdown (500 luxury SKUs) DQN ($K=9$) $37.5\%$ 15 days
Daily (1000 FMCGs) DQN ($K=100$) $5.10\times$ 30 days
Daily (1000 FMCGs) DDPG (continuous) $6.07\times$ 30 days
DDPG with continuous action space outperformed DQN with discrete bins across all daily pricing experiments, and both substantially outperformed manual expert pricing.